# Creation of an Fe3P Schreibersite Density Functional Tight Binding Model for Astrobiological Simulations

### v1

@@ 1 Introduction
Phosphorous plays a central role in the formation of many life-building compounds, including nucleotides, phospholipids, adenosine triphosphate (ATP) for metabolic energy storage, and calcium phosphate (apatite) found in bone.1, 2, 3, 4, 5 Extraterrestrial sources such as meteoritic impacts could have served as one possible source of prebiotic phosphorous on early Earth,4 including those with the mineral schreibersite, (Fe,Ni)3-P.6, 7 Schreibersite is a ferromagnetic material with tetrahedral symmetry (orthogonal lattice constants with $a=b>c$). It has been suggested to play a role in the formation of prebiotic organics,4, 8, 9, 6 where experimental studies have shown that it can catalyze the formation of amino acids, sugars, and nucleotides.1, 9, 10 However, early Earth likely experienced a wide range of thermodynamic conditions, including electrical storms11 and extreme thermodynamic conditions due to impacts12, 13, 14, that are difficult to probe with experimental trial and error alone. Schreibersite degradation studies would thus benefit from atomistic simulations, which can identify specific reaction mechanisms and kinetics, and can help narrow the relevant thermodynamic conditions for prebiotic synthesis.
To date, computational studies on schreibersite have largely focused on quantum simulations of bulk properties and water surface adsorption using Kohn-Sham Density Functional Theory (DFT).15, 16, 17 DFT is known to yield accurate descriptions of condensed phase physical and chemical data, as well as the energetics of chemical bond breaking and forming under reactive conditions (e.g., Refs. 12, 14). However, DFT simulations generally require heavy computational effort, which limits calculations to systems containing few hundreds of atoms and to molecular dynamics (MD) timescales of tens of picoseconds. These scales are generally orders of magnitude smaller than those probed by experiments, where for example, chemistry at elevated pressure and temperature can nanoseconds or longer to equilibrate,18, 19 and extended structures can form that are nanometer-scale in diameter.20, 21 Reactive MD models can help alleviate this time and length scale gap by optimizing empirical22 or machine-learning functional forms23, 24, 25, 26 to DFT or other quantum mechanical data. These methods are generally fit to the properties of specific chemical reactants over a narrow range of thermodynamic conditions. As a result, they can yield high computational efficiency, but tend to yield completely unphysical results outside of their fitting regime.
@@
In this work, we overcome these issues for Fe3P schreibersite by determining the Fe-Fe and Fe-P DFTB interaction potentials through a semi-automated workflow that uses the Chebyshev Interaction Model for Efficient Simulation (ChIMES) method for determination of the repulsive energies.45, 46, 47 ChIMES is a many-body reactive force field model for MD simulation based on linear combinations of Chebyshev polynomials. This linear parameterization yields rapid MD or $E_{\rm rep}$ model optimization, allowing us to efficiently cycle through large numbers of DFTB and ChIMES hyperparameters. DFTB/ChIMES models have been created for a wide variety of systems, including TiH2,48 energetic materials19, actinides49, and as a general purpose model for organic materials.50 Here, we first briefly review the DFTB and ChIMES approaches and explain our workflow for optimization of both DFTB and ChIMES hyperparameters. We then downselect our model through validation against a wide range of DFT data that was not included in our training set, such as the Fe3P lattice constants, bulk moduli, surface facet energies, and bulk water absorption. We then examine our optimal DFTB/ChIMES model further by computing surface facet energies, water surface adsorption on different facets, and MD simulations under ambient and extreme conditions. Finally, we explore the transferability of our model through the prediction of properties for different materials, including FeP, Fe2P, and three allotropes of iron. Our results indicate that our DFTB/ChIMES model yields comparable accuracy to DFT for all properties investigated here, including the intense pressures and temperatures that can be accessed during planetary impact.
@@ 2.1 DFT Calculations and Training Set Construction
  • 120 equally spaced configurations at $P=0$ GPa and $T=300,1500,$ and $2000$ K.
@@ 2.2 DFTB calculations
In general, optimal choice of confining radii is system-dependent and cannot be determined a priori, frequently requiring a semi-exhaustive search.67, 33 In our study, we choose to focus on optimizing the Fe density and wavefunction confining radii, only, in order to make our search more tractable. Similarly, we have only created $E_{\mathrm{Rep}}$ models for Fe-Fe and Fe-P interactions. The remaining DFTB$+$ interaction potentials (including tabulated SK Hamiltonian and overlap matrices, and/or $E_{\mathrm{Rep}}$ functions) were taken from known parameter sets. All non-iron interactions, such as those involving P, O, and H were taken from the mio-1-1 parameter set.27, 63 The Fe-O and Fe-H repulsive energies were determined by first computing new SK tables using our choice of confining radii for Fe and those from mio-1-1 for O and H. This was followed by matching energetics for dimer calculations from the trans3d parameter set44 to approximate the repulsive energies.
@@ 2.3 ChIMES Methodology and Many-body $E_{\mathrm{Rep}}$ determination
Here, the index $n_{\mathrm{a}}$ allows for a sum over the total number of atoms. The one-body contributions (${}^{1}\!E_{i_{1}}$) are element-specific constants, the two-body or dimer contributions (${}^{2}\!E_{i_{1}i_{2}}$) correspond to pair-wise interactions, the three-body or trimer contributions (${}^{3}\!E_{i_{1}i_{2}i_{3}}$) to triplet energies, etc., until a user-defined maximum of $n_{\mathrm{B}}$ is reached.
@@
Starting with the terms in the summation, $C_{m}^{e_{i_{1}}e_{i_{2}}}$ is the permutationally invariant coefficient for $m$th order polynomial (taken to maximum order of $N_{2B}$) for pairwise interactions between atoms $i$ and $j$, which are represented by element types $e_{i_{1}}$ and $e_{i_{2}}$. The term $s^{e_{i_{1}}e_{i_{2}}}_{i_{1}i_{2}}$ is the transformed pairwise distance ($r_{i_{1}i_{2}}$) such that the user-defined radial range $\left[r^{e_{i_{1}}e_{i_{2}}}_{\mathrm{min}}:r^{e_{i_{1}}e_{i_{2}}}_{\mathrm{max}}\right]$ is scaled to $\left[-1:1\right]$, using an element-pair specific scaling function. Similar to previous work45, we rely on the Morse-like exponential function, where $r_{i_{1}i_{2}}$ is scaled by a user chosen constant $\lambda_{e_{1}e_{2}}$, i.e., $s^{e_{i_{1}}e_{i_{2}}}_{i_{1}i_{2}}\propto\mathrm{exp}\left(-r_{i_{1}i_{2}}/\lambda_{e_{1}e_{2}}\right)$. The value of $\lambda_{e_{1}e_{2}}$ is usually taken to be the peak position of the first coordination shell from the radial distribution function for that element pair. We ensure smooth behavior and zero energy at the radial boundary ($r^{e_{i_{1}}e_{i_{2}}}_{\mathrm{max}}$) through the term $f^{e_{i_{1}}e_{i_{2}}}_{\mathrm{c}}(r_{i_{1}i_{2}})$, which we take to be is a Tersoff smoothing function68. Lastly, we use a smoothly varying penalty function $f_{\mathrm{p}}(r_{i_{1}i_{2}})$ in order to prevent sampling of distances below the minimum value sampled in our training set ($r^{e_{i_{1}}e_{i_{2}}}_{\mathrm{min}}$), thus preventing calculation of the Chebyshev polynomials outside of their natural range.
@@ 3.1 $R_{\Psi}$ and $R_{n}$ Optimization
Our results for the relative difference between the $\{8,4,0\}$ model and DFT for the schreibersite equilibrium lattice constants (Fig. 2 for the $a$ constant, with results for $c$ included in the Supporting Information for the sake of brevity) indicate that using equal weights in our optimizations for the forces and stress tensor components ($w_{\mathrm{F}}=w_{\rm\sigma}=1$) consistently yields an overly compact lattice, where the values of $a$ and $c$ are approximately $8\%$ and $10\%$ smaller than those from DFT. These higher density structures also optimize to antiferromagnetic spin geometries, rather than the correct ferromagnetic result. In addition, we observe a relatively small dependence on values of $R_{\Psi}$ and $R_{n}$. However, the introduction of weights of $w_{\rm\sigma}>1$ while keeping $w_{\mathrm{F}}$ at unity improves the results appreciably, where observe the errors improve to $-5\%$ for $w_{\rm\sigma}=5$ to $\sim 1\%$ for $w_{\rm\sigma}=25,50$.
Increasing the polynomial order to $\{16,12,0\}$ uniformly improves the lattice constant results relative to DFT, regardless of the value of $w_{\rm\sigma}$. Here, the average error in $a$ and $c$ is $\sim 3.5\%$ for $w_{\rm\sigma}=1$, which decreases to $<0.5\%$ for $w_{\rm\sigma}=25,50$. We note that we observe only small differences between results for $\{12,8,0\}$ and $\{12,8,4\}$, where the lattice constants both agree within 1% for values of $w_{\rm\sigma}=10$ and $25$ (see Supporting Information for detailed data).
@@
We now investigate the influence the confining radii on the schreibersite bulk modulus in order to test the validity of our models for high-pressure material property prediction. In this case, we have computed the zero temperature compression curves by isotropically compressing the system pressure to 100 GPa with each DFTB/ChIMES model. We then determine the bulk modulus $K_{0}$ through optimization to a Rose-Vinet equation of state.69 For reference, we previously obtained a DFT bulk modulus of $K_{\rm DFT}=173.2$ GPa15, compared to two available experimental results of 172.0172.0172.0172.070 and 157.9157.9157.9157.971 GPa.
@@
We now examine the interaction of schreibersite with water molecules in the bulk. We have used the PyMatGen code72 to identify two distinct interstitial sites within the lattice (Fig. 6). Specifically, the first type of interstitial site (site 1) consists of a tetrahedral cage composed of iron atoms. The second type of interstitial site (site 2) is formed by a tetrahedron that incorporates a phosphorus atom at one of its vertices. The formation energies for these sites are calculated via the equation $E_{\rm def}=E_{\rm Fe_{3}P+H_{2}O}-(E_{\rm bulk}+E_{\rm H_{2}O})$, where $E_{\rm Fe_{3}P+H_{2}O}$ represents the total energy of the relaxed defected structure, $E_{\rm bulk}$ is the total energy of the bulk schreibersite, and $E_{\rm H_{2}O}$ denotes the energy of an isolated water molecule in vacuum. In this case, both DFT and all DFTB/ChIMES calculations were initiated from the same initial configurations, which resulted in dissociation of the water molecule upon optimization. For reference, the values computed via DFT are $E_{\rm site1}=-3.56$ eV and $E_{\rm site2}=-3.49$ eV, respectively.
@@ 3.2 Water Surface Absorption Energies
Figure 8: Different adsorption sites on $(110)$ surface. In this image, the big brown circles represent iron atoms, while the smaller orange circles stand for the phosphorous atoms. Graphic taken from Ref. 15.
@@
In contrast, the adsorption energies from DFTB/ChIMES predominantly involving phosphorus atoms tend to be more accurate, where the Fe-P bridge site results differ from DFT by less than 0.05 eV and the P ontop site by less than 0.04 eV. These discrepancies in the Fe results are likely due in part to the fact that the Fe-O and Fe-H pairwise repulsive energies were determined by matching to the trans3d parameter set, which was optimized for periodic solids, whereas inclusion of chemical surface interactions can require additional optimizations.67 However, the mio-1-1 parameters used for O, H, and P interactions were optimized to a properties of a relatively large series different molecules, yielding improved surface-molecule interactions.
@@ 3.3 Magentization as a Function of Pressure
Given one of our objectives of developing a model suitable for high-pressure impact studies, we consider the variation of magnetization in response to external pressure. Here, we have computed the relative magnetization in scheibersite as it is compressed at zero Kelvin from zero to 100 GPa (Fig. 9). Results from DFTB/ChIMES and DFT are very similar throughout the entire pressure range, with a gradual decrease up to 30 GPa, where we observe a rapid drop that corresponds to a potential first-order phase transition.71, 70, 73 This diminishing trend culminates in a non-magnetic state at $P>50$ GPa in both DFT and DFTB/ChIMES, where the iron atoms have largely lost their magnetic moments, as previously noted by high-pressure experiments.73

### v2

@@ 1 Introduction
Phosphorous plays a central role in the formation of many life-building compounds, including nucleotides, phospholipids, adenosine triphosphate (ATP) for metabolic energy storage, and calcium phosphate (apatite) found in bone.1, 2, 3, 4, 5 Extraterrestrial sources such as meteoritic impacts could have served as one possible source of prebiotic phosphorous on early Earth,4 including those with the mineral schreibersite, (Fe,Ni)3-P.6, 7 Schreibersite is a ferromagnetic material with tetrahedral symmetry (orthogonal lattice constants with $a=b>c$). It has been suggested to play a role in the formation of prebiotic organics,4, 8, 9, 6 where experimental studies have shown that it can catalyze the formation of amino acids, sugars, and nucleotides.1, 9, 10 However, early Earth likely experienced a wide range of conditions, including electrical storms11 and extreme thermodynamic conditions due to impacts12, 13, 14, that are difficult to probe with experimental trial and error alone. Schreibersite degradation studies would thus benefit from atomistic simulations, which can identify specific reaction mechanisms and kinetics, and can help narrow the relevant thermodynamic conditions for prebiotic synthesis.
To date, computational studies on schreibersite have largely focused on quantum simulations of bulk properties and water surface adsorption using Kohn-Sham Density Functional Theory (DFT).15, 16, 17 DFT is known to yield accurate descriptions of condensed phase physical and chemical data, as well as the energetics of chemical bond breaking and forming under reactive conditions (e.g., Refs. 12, 14). However, DFT simulations generally require heavy computational effort, which limits calculations to systems containing few hundreds of atoms and to molecular dynamics (MD) timescales of tens of picoseconds. These scales are generally orders of magnitude smaller than those probed by experiments, where for example, chemistry at elevated pressure and temperature can require nanoseconds or longer to equilibrate,18, 19 and extended structures can form that are nanometer-scale in diameter.20, 21 Reactive MD models can help alleviate this time and length scale gap by optimizing empirical22 or machine-learning functional forms23, 24, 25, 26 to DFT or other quantum mechanical data. These methods are generally fit to the properties of specific chemical reactants over a narrow range of thermodynamic conditions. As a result, they can yield high computational efficiency, but tend to yield completely unphysical results outside of their fitting regime.
@@
In this work, we overcome these issues for Fe3P schreibersite by determining the Fe-Fe and Fe-P DFTB interaction potentials through a semi-automated workflow that uses the Chebyshev Interaction Model for Efficient Simulation (ChIMES) method for determination of the repulsive energies.45, 46, 47 ChIMES is a many-body reactive force field model for MD simulation based on linear combinations of Chebyshev polynomials. This linear parameterization yields rapid MD or $E_{\rm rep}$ model optimization, allowing us to efficiently cycle through large numbers of DFTB and ChIMES hyperparameters. DFTB/ChIMES models have been created for a wide variety of systems, including TiH2,48 energetic materials19, actinides49, and as a general purpose model for organic materials.50 Here, we first briefly review the DFTB and ChIMES approaches and explain our workflow for optimization of both DFTB and ChIMES hyperparameters. We then downselect our model through validation against a wide range of DFT data that was not included in our training set, such as the Fe3P lattice constants, bulk moduli, surface facet energies, and bulk water absorption. We then examine our optimal DFTB/ChIMES model further by computing water surface adsorption energies, the phonon density of states of the perfect crystal, and MD simulations under ambient and extreme conditions. Finally, we explore the transferability of our model through the prediction of properties for different materials, including FeP, Fe2P, and three allotropes of iron. Our results indicate that our DFTB/ChIMES model yields comparable accuracy to DFT for all properties investigated here, including the intense pressures and temperatures that can be accessed during planetary impact.
@@ 2.1 DFT Calculations and Training Set Construction
  • 120 configurations equally spaced in time (e.g., sampled every $\sim$167 fs from the MD trajectory) at $P=0$ GPa and $T=300,1500,$ and $2000$ K.
@@ 2.2 DFTB calculations
In general, optimal choice of confining radii is system-dependent and cannot be determined a priori, frequently requiring a semi-exhaustive search.67, 33 In our study, we choose to focus on optimizing the Fe density and wavefunction confining radii, only, in order to make our search more tractable. Similarly, we have only created $E_{\mathrm{Rep}}$ models for Fe-Fe and Fe-P interactions. The remaining DFTB$+$ interaction potentials (including tabulated SK Hamiltonian and overlap matrices, and/or $E_{\mathrm{Rep}}$ functions) were taken from known parameter sets. All non-iron interactions, such as those involving P, O, and H were taken from the mio-1-1 parameter set.27, 63 The Fe-O and Fe-H repulsive energies were determined by first computing new SK tables using our choice of confining radii for Fe and those from mio-1-1 for O and H. This was followed by matching energetics for dimer calculations from the trans3d parameter set44 to approximate the repulsive energies. A recent publication has made available DFTB parameterizations for a large portion of the periodic table68, though these were not available at the time of this study and have not been tested for schreibersite/water systems, specifically.
@@ 2.3 ChIMES Methodology and Many-body $E_{\mathrm{Rep}}$ determination
Here, the index $n_{\mathrm{a}}$ allows for a sum over the total number of atoms. The one-body contributions (${}^{1}\!E_{i_{1}}$) are element-specific constants, the two-body (2B) or dimer contributions (${}^{2}\!E_{i_{1}i_{2}}$) correspond to pair-wise interactions, the three-body (3B) or trimer contributions (${}^{3}\!E_{i_{1}i_{2}i_{3}}$) to triplet energies, etc., until a user-defined maximum of $n_{\mathrm{B}}$ is reached.
@@
Starting with the terms in the summation, $C_{m}^{e_{i_{1}}e_{i_{2}}}$ is the permutationally invariant coefficient for $m$th order polynomial (taken to maximum order of $N_{2B}$) for pairwise interactions between atoms $i$ and $j$, which are represented by element types $e_{i_{1}}$ and $e_{i_{2}}$. The term $s^{e_{i_{1}}e_{i_{2}}}_{i_{1}i_{2}}$ is the transformed pairwise distance ($r_{i_{1}i_{2}}$) such that the user-defined radial range $\left[r^{e_{i_{1}}e_{i_{2}}}_{\mathrm{min}}:r^{e_{i_{1}}e_{i_{2}}}_{\mathrm{max}}\right]$ is scaled to $\left[-1:1\right]$, using an element-pair specific scaling function. Similar to previous work45, we rely on the Morse-like exponential function, where $r_{i_{1}i_{2}}$ is scaled by a user chosen constant $\lambda_{e_{1}e_{2}}$, i.e., $s^{e_{i_{1}}e_{i_{2}}}_{i_{1}i_{2}}\propto\mathrm{exp}\left(-r_{i_{1}i_{2}}/\lambda_{e_{1}e_{2}}\right)$. The value of $\lambda_{e_{1}e_{2}}$ is usually taken to be the peak position of the first coordination shell from the radial distribution function for that element pair. We ensure smooth behavior and zero energy at the radial boundary ($r^{e_{i_{1}}e_{i_{2}}}_{\mathrm{max}}$) through the term $f^{e_{i_{1}}e_{i_{2}}}_{\mathrm{c}}(r_{i_{1}i_{2}})$, which we take to be is a Tersoff smoothing function69. Lastly, we use a smoothly varying penalty function $f_{\mathrm{p}}(r_{i_{1}i_{2}})$ in order to prevent sampling of distances below the minimum value sampled in our training set ($r^{e_{i_{1}}e_{i_{2}}}_{\mathrm{min}}$), thus preventing calculation of the Chebyshev polynomials outside of their natural range.
@@ 3.1 $R_{\Psi}$ and $R_{n}$ Optimization
Our results for the relative difference between the $\{8,4,0\}$ model and DFT for the schreibersite equilibrium lattice constants (Fig. 2 for the $a$ constant, with results for $c$ included in the Supporting Information) indicate that using equal weights in our optimizations for the forces and stress tensor components ($w_{\mathrm{F}}=w_{\rm\sigma}=1$) consistently yields an overly compact lattice, where the values of $a$ and $c$ are approximately $8\%$ and $10\%$ smaller than those from DFT. These higher density structures also optimize to antiferromagnetic spin geometries, rather than the correct ferromagnetic result. In addition, we observe a relatively small dependence on values of $R_{\Psi}$ and $R_{n}$. However, the introduction of weights of $w_{\rm\sigma}>1$ while keeping $w_{\mathrm{F}}$ at unity improves the results appreciably, where observe the errors improve to $-5\%$ for $w_{\rm\sigma}=5$ to $\sim 1\%$ for $w_{\rm\sigma}=25,50$.
Increasing the polynomial order to $\{16,12,0\}$ uniformly improves the lattice constant results relative to DFT. Here, the average error in $a$ and $c$ is $\sim 3.5\%$ for $w_{\rm\sigma}=1$, which decreases to $<0.5\%$ for $w_{\rm\sigma}=25,50$. We observe only small differences between results for $\{12,8,0\}$ and $\{12,8,4\}$, where the lattice constants both agree within 1% for values of $w_{\rm\sigma}=10$ and $25$ (see Supporting Information for detailed data).
@@
We now investigate the influence the confining radii on the schreibersite bulk modulus in order to test the validity of our models for high-pressure material property prediction. In this case, we have computed the zero temperature compression curves by isotropically compressing the system pressure to 100 GPa with each DFTB/ChIMES model. We then determine the bulk modulus $K_{0}$ through optimization to a Rose-Vinet equation of state.70 For reference, we previously obtained a DFT bulk modulus of $K_{\rm DFT}=173.2$ GPa15, compared to two available experimental results of 172.0172.0172.0172.071 and 157.9157.9157.9157.972 GPa.
@@
We now examine the interaction of schreibersite with water molecules in the bulk. We have used the PyMatGen code73 to identify two distinct interstitial sites within the lattice (Fig. 6). Specifically, the first type of interstitial site (site 1) consists of a tetrahedral cage composed of iron atoms. The second type of interstitial site (site 2) is formed by a tetrahedron that incorporates a phosphorus atom at one of its vertices. The formation energies for these sites are calculated via the equation $E_{\rm def}=E_{\rm Fe_{3}P+H_{2}O}-(E_{\rm bulk}+E_{\rm H_{2}O})$, where $E_{\rm Fe_{3}P+H_{2}O}$ represents the total energy of the relaxed defected structure, $E_{\rm bulk}$ is the total energy of the bulk schreibersite, and $E_{\rm H_{2}O}$ denotes the energy of an isolated water molecule in vacuum. In this case, both DFT and all DFTB/ChIMES calculations were initiated from the same initial configurations, which resulted in dissociation of the water molecule upon optimization. For reference, the values computed via DFT are $E_{\rm site1}=-3.56$ eV and $E_{\rm site2}=-3.49$ eV, respectively.
@@ 3.2 Water Surface Absorption Energies
Figure 8: Different adsorption sites on $(110)$ surface. In this image, the big brown circles represent iron atoms, while the smaller orange circles stand for the phosphorous atoms. Reproduced from Ref. 15. Copyright 2022 American Chemical Society.
@@
In contrast, the adsorption energies from DFTB/ChIMES predominantly involving phosphorus atoms tend to be more accurate, where the Fe-P bridge site results differ from DFT by less than 0.05 eV and the P ontop site by less than 0.04 eV. These discrepancies in the Fe results are likely due in part to the fact that the Fe-O and Fe-H pairwise repulsive energies were determined by matching to the trans3d parameter set, which were optimized for molecular data, whereas inclusion of periodic solids as well as chemical surface interactions can require additional optimizations.67 However, the mio-1-1 parameters used for O, H, and P interactions were optimized to a properties of a relatively large series different molecules, yielding improved surface-molecule interactions.
@@ 3.3 Magentization as a Function of Pressure
Given one of our objectives of developing a model suitable for high-pressure impact studies, we consider the variation of magnetization in response to external pressure. Here, we have computed the relative magnetization in scheibersite as it is compressed at zero Kelvin from zero to 100 GPa (Fig. 9). Results from DFTB/ChIMES and DFT are very similar throughout the entire pressure range, with a gradual decrease up to 30 GPa, where we observe a rapid drop that corresponds to a potential first-order phase transition.72, 71, 74 This diminishing trend culminates in a non-magnetic state at $P>50$ GPa in both DFT and DFTB/ChIMES, where the iron atoms have largely lost their magnetic moments, as previously noted by high-pressure experiments.74
@@
3.4 Phonon Density of States
Figure 10: Phonon density of states, with results from DFTB/ChIMES in red and those from DFT in black.
We now examine the accuracy of the vibrational frequency from DFTB/ChIMES by computing the total phonon density of states (Fig. 10). These were computed from a zero-temperature optimized primitive cell with the phonopy code,75, 76 using a $2\times 2\times 2$ supercell and a $31\times 31\times 31$ q-point (wavevector) mesh. Overall, we observe reasonable agreement between DFTB/ChIMES and DFT. The manifold of vibrational states between zero and $\sim$300 cm-1 shows close agreement up to 250 cm-1, though DFTB/ChIMES shows an enhanced density of states between 250-300 cm-1 that is not present in the DFT results. Both DFT and DFTB/ChIMES exhibit a doublet above 300 cm-1, though the DFTB/ChIMES result occurs $\sim$50 cm-1 higher. The remaining manifold of peaks from DFT occurs roughly between 360–430 cm-1, and is somewhat different in shape compared to the DFTB/ChIMES result, which spans approximately 390–445 cm-1. These deviations are relatively in magnitude, given that a difference of 50 cm-1, for example, corresponds to a change in energy of $\sim 6\times 10^{-4}$ eV.